OpenStudy (fibonaccichick666):

Quick help please: Prove that * defined on the rationals is a group by letting a*b=ab.

OpenStudy (fibonaccichick666):

I have group axioms 1,2 and closed covered, but for some reason I can't figure out how to show the inverse in the rationals.

OpenStudy (fibonaccichick666):

@amistre64 , could you take a quick look?

OpenStudy (amistre64):

inverses are defined as: x x' = e

OpenStudy (amistre64):

group properties (in my text) were defined as: identity, inverses, associativity; for a closed binary operation

OpenStudy (fibonaccichick666):

right ooh ok, so let's say a=c/d and b=f/g so ab=(cf)/(dg) which is an element of the rationals, so (ab)^-1=(dg)/(cf) right?

OpenStudy (fibonaccichick666):

yea, it's just inverses I seem to always get stuck on proving but, same

OpenStudy (amistre64):

yes

OpenStudy (fibonaccichick666):

thank you!!

OpenStudy (amistre64):

let a = n/m for some integers m,n; m not 0 \[\frac nma'=e\] \[\frac{m}{1}\frac nma'=\frac{m}{1}\] \[\frac n1a'=\frac{m}{1}\] \[\frac 1n\frac n1a'=\frac{m}{1}\frac{1}{n}\] \[a'=\frac{m}{n}\in Q\]

OpenStudy (amistre64):

or some such reasoning :) yours was fine for me

OpenStudy (fibonaccichick666):

thanks, ya know sometimes you just have to rubber duck with someone haha thanks again!